OpenStudy (studybird123):

Write the ratio in three different formats: to , , and . Carlos biked 24 miles and ran 6 miles. Write the ratio of the distance Carlos biked to the total distance. A. 6 to 24, 6:24, 6/24 B. 3 to 12, 3:12, 3/12 C. 30 to 24, 30:24, 30/24 D. 4 to 5, 4:5, 4/5

OpenStudy (studybird123):

Is it d?

OpenStudy (acxbox22):

biked= 24 total distance= 24+6 so it would be 24/30=4/5 you are correct

OpenStudy (studybird123):

oh thank you!!